anne had 1 pizza, eriston had 1 3/4 pizzas and julie had 1/2 of a pizza. find the total number of pizzas

Answer by greenestamps(13200)



Picture the actual pizzas....

When you cut a whole pizza into two equal size pieces, you get two half pizzas. Cut each of those two halves into two equal size pieces, you get four equal size pizzas; each piece is one quarter of a pizza.

So to add the given numbers of pizzas, cut everything into the same size pieces -- quarters of a pizza.

1 pizza --> 4 quarters
1 3/4 pizzas --> the whole is 4 quarters; plus 3 more quarters, so 7 quarters
1/2 pizza --> 2 quarters

Now all the pieces are the same size, so find how many there are all together: 4+7+2 = 13 quarters.

So one form of the answer to the question is 13/4 pizzas.

Then you can think of putting the quarters together to make whole pizzas. 13 = 3(4)+1; so with the 13 quarter pizzas you can make 3 whole pizzas with one quarter left over.

So another form of the answer is 3 1/4 pizzas.

Working with the numbers themselves instead of the actual pizzas, the process is exactly the same. To add the fractions, you need a common denominator (same size pieces). Since the least common denominator of 2 and 4 is 4,

1 = 1*(4/4) = 4/4
1 3/4 = 1*(4/4) + 3/4 = 4/4 + 3/4 = 7/4
1/2 = (1/2)*(2/2) = 2/4

Then (adding same size pieces)

4/4 + 7/4 + 2/4 = 13/4

Finally, if you need an answer as a mixed number instead of an improper fraction, you can put some of the pieces together to make wholes, with some left over:

13/4 = 3 remainder 1 --> 3 1/4 (3 whole pizzas with one quarter left over)